# Classification of nilpotent almost abelian Lie groups admitting left-invariant complex or symplectic structures

### v1

@@ 4. Classification of nilpotent almost abelian Lie algebras admitting symplectic structures
and it turns out that each $\mathfrak{sp}_{\Omega}(2m,{\mathbb{R}})$ is conjugate to $\mathfrak{sp}(2m,{\mathbb{R}})$ in $GL(2m,{\mathbb{R}})$.
@@
The next proposition gives a characterization of the $2n$-dimensional almost abelian Lie algebras admitting a symplectic structure (see also [30, Proposition 4.1]).
@@
where the last equality holds from $\omega(x,e_{1})=0$ for all $x\in\mathfrak{h}$. Hence, $\omega_{\mathfrak{u}}(v,z)=0$ for all $z\in\mathfrak{u}$, which implies that $v=0$ since $\omega_{\mathfrak{u}}$ is non-degenerate.
Finally, since $\omega$ is closed, $\mathfrak{u}$ is abelian and $\omega(x,e_{1})=0$ for all $x\in\mathfrak{u}$, it follows that $\Pi\circ\left(\text{ad}_{e_{0}}|_{\mathfrak{u}}\right)\in\mathfrak{sp}(\mathfrak{u},\omega_{\mathfrak{u}})$. ∎
@@
The following fact will be used frequently in the sequel.
Remark 4.4.
The next well-known lemma (see for instance, [14, Theorem 5.1.3]) will be useful to prove existence of symplectic structures on certain almost abelian Lie algebras.
@@
The matrix $\mathcal{J}_{2k}$ as in (1) is conjugate to a matrix in $\mathfrak{sp}(2k,{\mathbb{R}})$ for all $k\geq 1$.
@@
For $r=2k-1$, the terms in the expression of $B_{2k-1}$ corresponding to $i\neq k$ vanish, that is, $B_{2k-1}=(-1)^{k-1}X^{k-1}(X^{T})^{k-1}$, which is the matrix whose $(k,k)$ coefficient is $(-1)^{k-1}$ and the remaining entries are equal to zero. In particular, $B_{2k-1}\neq 0$, which implies that $N^{2k-1}\neq 0$. Therefore, $N$ is conjugate to $\mathcal{J}_{2k}$, and the lemma follows. ∎
@@
Throughout the proof, we will use several times Remark 4.4 and the facts that $(\mathcal{J}_{3}\oplus-\mathcal{J}_{3}^{T})\in\mathfrak{sp}\left(6,{\mathbb{R}}\right)$, $\mathcal{J}_{2}\in\mathfrak{sp}\left(2,{\mathbb{R}}\right)$, $0_{2k}\in\mathfrak{sp}\left(2k,{\mathbb{R}}\right)$ (see (9)) and $\mathcal{J}_{4}\in\mathfrak{sp}(4,{\mathbb{R}})$ (by Lemma  4.5).
@@
We exhibit next an example of a 5-step nilpotent almost abelian Lie algebra which does not admit a symplectic structure.
@@
Let $\mathfrak{g}_{M}={\mathbb{R}}e_{0}\ltimes_{M}{\mathbb{R}}^{9}$ be the $10$-dimensional almost abelian Lie algebra with $M=\mathcal{J}_{5}\oplus\mathcal{J}_{3}\oplus 0_{1}$. We claim that $\mathfrak{g}_{M}$ does not admit a symplectic structure. Indeed, let =$\mathcal{B}=\{e_{0},\dots,e_{9}\}$ be the basis of $\mathfrak{g}_{M}$ such that $M$ is the matrix of $\operatorname{ad}_{e_{0}}|_{{\mathbb{R}}^{9}}$ with respect to $\{e_{1},\dots,e_{9}\}$ and denote by $\{e^{0},\dots,e^{9}\}$ the dual basis of $\mathcal{B}$. We have that,
@@
Hence, any closed 2-form on $\mathfrak{g}$ can be written as $\omega=\sum_{j=1}^{9}a_{j}e^{0j}+\beta$ where $\beta$ is given by:
@@
4.1. Classification Theorem (Symplectic Case)
Theorem 4.10 below gives the classification of the nilpotent almost abelian Lie algebras admitting a symplectic structure. For its proof we will need some facts.
Combining [14, Theorem 5.1.3] with [26, §1.4] the next theorem follows.
@@ 4.1. Classification Theorem (Symplectic Case)
Let $D=\mathcal{J}_{m_{1}}^{\oplus q_{1}}\oplus\cdots\oplus\mathcal{J}_{m_{r}}^{\oplus q_{r}}\oplus 0_{s}$ with $m_{1}>\cdots>m_{r}\geq 2$, $q_{i}\geq 1$ for all $i$, $s\geq 0$ and $\sum_{i=1}^{r}q_{i}m_{i}+s=2m$. Then, $D$ is conjugate in $GL(2m,{\mathbb{R}})$ to a matrix in $\mathfrak{sp}(2m,{\mathbb{R}})$ if and only if $s$ is even and $q_{i}$ is even for all $i$ such that $m_{i}$ is odd.
Lemma 4.9.
Let $\mathfrak{g}_{A}$ be a nilpotent almost abelian Lie algebra, $\dim\mathfrak{g}_{A}=2(m+1)$. Then $\mathfrak{g}_{A}$ admits a symplectic structure if and only if $A$ is conjugate to
$C=\begin{pmatrix}0&0\\ w&D\end{pmatrix},\quad\text{where }w\in\mathbb{R}^{2m},\;D\in\mathfrak{sp}(2m,{\mathbb{R}}).$
Proof.
$C=\begin{pmatrix}0&v^{T}\\ 0&B\end{pmatrix},\quad\text{where }v\in\mathbb{R}^{2m},\;B\in\mathfrak{sp}_{\Omega}(2m,{\mathbb{R}}),$
for certain non-singular skew-symmetric matrix $\Omega$. It follows from Remark 4.1 that $C$ is conjugate to
$C^{\prime}=\begin{pmatrix}0&(v^{\prime})^{T}\\ 0&B^{\prime}\end{pmatrix},\quad\text{where }v^{\prime}\in\mathbb{R}^{2m},\;B^{\prime}\in\mathfrak{sp}(2m,{\mathbb{R}}).$
Since $C^{\prime}$ is conjugate to its transpose and $(B^{\prime})^{T}\in\mathfrak{sp}(2m,{\mathbb{R}})$, then $A$ is conjugate to
$\begin{pmatrix}0&0\\ v^{\prime}&(B^{\prime})^{T}\end{pmatrix},\quad\text{where }v^{\prime}\in\mathbb{R}^{2m},\;(B^{\prime})^{T}\in\mathfrak{sp}(2m,{\mathbb{R}}),$
and the lemma follows with $w=v^{\prime}$ and $D=(B^{\prime})^{T}.$
Theorem 4.10.
Let $\mathfrak{g}_{M}={\mathbb{R}}e_{0}\ltimes_{M}{\mathbb{R}}^{2n-1}$ be the nilpotent almost abelian Lie algebra with
$M=\mathcal{J}_{n_{1}}^{\oplus p_{1}}\oplus\cdots\oplus\mathcal{J}_{n_{k}}^{\oplus p_{k}}\oplus 0_{t},\quad n_{1}>\cdots>n_{k}\geq 2,\;\,p_{i}\geq 1\text{for all }i,\;t\geq 0,$
where $\sum_{i=1}^{k}p_{i}n_{i}+t=2n-1$. Then $\mathfrak{g}_{M}$ admits a symplectic structure if and only if one of the following conditions holds:
@@ 2. ${\rm(ii)}$
We suppose that $\mathfrak{g}_{M}$ admits a symplectic structure. Hence, by Lemma 4.9, $M$ is conjugate to $C=\begin{pmatrix}0&0\\ w&D\end{pmatrix}$, where $w\in\mathbb{R}^{2n-2}$ and $D\in\mathfrak{sp}(2(n-1),{\mathbb{R}})$. Since $D$ is a nilpotent matrix, Theorem 4.8 implies that $D$ is conjugate to
@@
Next we consider that $M$ is conjugate to case (b). In this case $t$ is even, since $t=s$. We suppose that $m_{d}$ is even, so $m_{d}+1$ is odd and $m_{d-1}\geq m_{d}+1$. If $m_{d-1}>m_{d}+1$, then $\mathcal{J}_{m_{d}+1}$ appears only once. If $m_{d-1}=m_{d}+1$, then $m_{d-1}$ is odd. So Theorem 4.8 implies that $q_{d-1}$ is even. In this way, $\mathcal{J}_{m_{d}+1}$ appears an odd number of times. Now, we suppose that $m_{d}$ is odd. By Theorem 4.8, $q_{d}$ is even and Corollary 2.6 says that $\mathcal{J}_{m_{d}}$ appears $q_{d}-1$ times, that is, an odd number of times. Finally, the uniqueness in condition (ii) follows from Theorem 4.8.
@@
$t-1$ is even and $p_{i}$ is even for all $i$ such that $n_{i}$ is odd. Applying Theorem 4.8, $D$ is conjugate to a matrix in $\mathfrak{sp}(2(n-1),{\mathbb{R}})$. By Remark 4.3, $\mathfrak{g}_{M}$ admits a symplectic structure.
@@
As a consequence of Theorems 3.10 and 4.10 we obtain the next result, whose proof is straightforward.
@@
The above theorem does not hold for arbitrary nilpotent Lie algebras, as the next example shows.

### v2

@@ 5.1. Characterization of almost abelian Lie algebras admitting a symplectic structure
and it turns out that each $\mathfrak{sp}_{\Omega}(2m,{\mathbb{R}})$ is conjugate to $\mathfrak{sp}(2m,{\mathbb{R}})$ in $GL(2m,{\mathbb{R}})$. More precisely, if $\Omega=Q^{T}J_{m}Q$ for some $Q\in GL(2m,{\mathbb{R}})$, then $D\in\mathfrak{sp}_{\Omega}(2m,{\mathbb{R}})$ if and only if $QDQ^{-1}\in\mathfrak{sp}(2m,{\mathbb{R}})$. Moreover, since $D^{T}\in\mathfrak{sp}(2m,{\mathbb{R}})$ for all $D\in\mathfrak{sp}(2m,{\mathbb{R}})$, it follows that $B\in\mathfrak{sp}_{\Omega_{1}}(2m,{\mathbb{R}})$ if and only if $B^{T}\in\mathfrak{sp}_{\Omega_{2}}(2m,{\mathbb{R}})$, where $\Omega_{1}=P^{T}J_{m}P$ and $\Omega_{2}=P^{-1}J_{m}(P^{-1})^{T}$, for some $P\in GL(2m,{\mathbb{R}})$.
@@
The following fact will be used frequently in the sequel.
Remark 5.2.
The next proposition gives a characterization of the $2n$-dimensional almost abelian Lie algebras admitting a symplectic structure (see also [33, Proposition 4.1]).
@@
where the last equality holds from $\omega(x,e_{1})=0$ for all $x\in\mathfrak{h}$. Hence, $\omega_{\mathfrak{u}}(v,z)=0$ for all $z\in\mathfrak{u}$, which implies that $v=0$ since $\omega_{\mathfrak{u}}$ is non-degenerate. Finally, since $\omega$ is closed, $\mathfrak{u}$ is abelian and $\omega(x,e_{1})=0$, for all $x\in\mathfrak{u}$, and it follows that $\Pi\circ\left(\text{ad}_{e_{0}}|_{\mathfrak{u}}\right)\in\mathfrak{sp}(\mathfrak{u},\omega_{\mathfrak{u}})$. ∎
@@
Corollary 5.5.
Let $\mathfrak{g}_{A}$ be a $2m$-dimensional almost abelian Lie algebra. Then $\mathfrak{g}_{A}$ admits a symplectic structure if and only if $A$ is conjugate to
(17) $C=\begin{pmatrix}c&0\\ w&D\end{pmatrix},\quad\text{where }c\in{\mathbb{R}},\,w\in\mathbb{R}^{2(m-1)},\;D\in\mathfrak{sp}_{\Omega}(2(m-1),{\mathbb{R}}),$
for some non-singular skew-symmetric $\Omega\in M(2(m-1),\mathbb{R})$.
Proof.
$C=\begin{pmatrix}c&v^{T}\\ 0&B\end{pmatrix},\quad\text{where }c\in{\mathbb{R}},\,v\in\mathbb{R}^{2(m-1)},\;B\in\mathfrak{sp}_{\Omega_{1}}(2(m-1),{\mathbb{R}}),$
with $\Omega_{1}=P^{T}J_{m}P$, for some $P\in GL(2(m-1),{\mathbb{R}})$. Since $C$ is conjugate to its transpose, then $A$ is conjugate to $\begin{pmatrix}c&0\\ v&B^{T}\end{pmatrix}$ where $B^{T}\in\mathfrak{sp}_{\Omega_{2}}(2(m-1),{\mathbb{R}})$ with $\Omega_{2}=P^{-1}J_{m}(P^{-1})^{T}$ (see Remark 5.1). The corollary follows by taking $\Omega=\Omega_{2}$. ∎
There is a classical theorem describing nilpotent orbits of the Lie algebra $\mathfrak{sp}(2m,{\mathbb{R}})$ (see [15, Chapter 5]). This result was generalized in [35, Theorem 2.4] to any orbit of $\mathfrak{sp}(2m,{\mathbb{R}})$. The next theorem, which will be useful for our purposes, is a consequence of [35, Theorem 2.4].
Theorem 5.6.
Let $\Omega\in M(2m,{\mathbb{R}})$ be a non-singular skew-symmetric matrix and let $B\in M(2m,{\mathbb{R}})$. Then $B$ is conjugate in $GL(2m,{\mathbb{R}})$ to a matrix in $\mathfrak{sp}_{\Omega}(2m,{\mathbb{R}})$ if and only if any $\lambda\in\emph{spec}\,(B)$ satisfies:
(18) $\text{if }\lambda=0\text{ and }\Lambda(B(0))=(n_{1},\ldots,n_{k};p_{1},\ldots,p_{k};t),\text{ then }p_{r}\text{ is even for all odd }n_{r}\text{ and }t\text{ is}$
[…]
(19) $\text{if }\lambda\notin i{\mathbb{R}},\text{ then }-\lambda\in\emph{spec}\,(B)\text{ and }\Lambda(B(-\lambda))=\Lambda(B(\lambda)).$
In particular, if $B$ is conjugate to a matrix in $\mathfrak{sp}_{\Omega}(2m,{\mathbb{R}})$ then its semisimple and nilpotent parts are conjugate to matrices in $\mathfrak{sp}_{\Omega_{1}}(2m,{\mathbb{R}})$ and $\mathfrak{sp}_{\Omega_{2}}(2m,{\mathbb{R}})$, respectively, for some $\Omega_{1}$ and $\Omega_{2}$.
Proof.
The first part of the theorem follows from [35, Theorem 2.4].
To prove the last assertion, decompose $B$ as $B=B_{s}+B_{n}$ with $B_{s}$ semisimple and $B_{n}$ nilpotent. It is easily verified that if $B$ satisfies conditions (18) and (19), then so do $B_{n}$ and $B_{s}$, hence the last assertion follows. ∎
5.2. The nilpotent case
The next well-known lemma (see for instance, [15, Theorem 5.1.3]) will be useful to prove existence of symplectic structures on certain almost abelian Lie algebras.
@@ 5.2. The nilpotent case
The matrix $\mathcal{J}_{2k}$ is conjugate to a matrix in $\mathfrak{sp}(2k,{\mathbb{R}})$ for all $k\geq 1$.
@@
For $r=2k-1$, the terms in the expression of $B_{2k-1}$ corresponding to $i\neq k$ vanish. Hence, $B_{2k-1}=(-1)^{k-1}X^{k-1}(X^{T})^{k-1}$, which is the matrix whose $(k,k)$ coefficient is $(-1)^{k-1}$ and the remaining entries are equal to zero. In particular, $B_{2k-1}\neq 0$, which implies that $N^{2k-1}\neq 0$. Therefore, $N$ is conjugate to $\mathcal{J}_{2k}$, and the lemma follows. ∎
@@
Throughout the proof, we will use several times Remark 5.2 and the facts that $(\mathcal{J}_{3}\oplus-\mathcal{J}_{3}^{T})\in\mathfrak{sp}\left(6,{\mathbb{R}}\right)$, $\mathcal{J}_{2}\in\mathfrak{sp}\left(2,{\mathbb{R}}\right)$, $0_{2k}\in\mathfrak{sp}\left(2k,{\mathbb{R}}\right)$ (see (14)) and $\mathcal{J}_{4}\in\mathfrak{sp}(4,{\mathbb{R}})$ (by Lemma  5.7).
@@
We show below an example of a 5-step nilpotent almost abelian Lie algebra which does not admit a symplectic structure.
@@
Let $\mathfrak{g}_{M}={\mathbb{R}}e_{0}\ltimes_{M}{\mathbb{R}}^{9}$ be the $10$-dimensional almost abelian Lie algebra with $M=\mathcal{J}_{5}\oplus\mathcal{J}_{3}\oplus 0_{1}$. We claim that $\mathfrak{g}_{M}$ does not admit a symplectic structure. Indeed, let =$\mathcal{B}=\{e_{0},\dots,e_{9}\}$ be the basis of $\mathfrak{g}_{M}$ such that $M$ is the matrix of $\operatorname{ad}_{e_{0}}|_{{\mathbb{R}}^{9}}$ with respect to $\{e_{1},\dots,e_{9}\}$ and denote by $\{e^{0},\dots,e^{9}\}$ the dual basis of $\mathcal{B}$.
We have that,
@@
Hence, any closed 2-form on $\mathfrak{g}$ can be written as $\omega=\sum_{j=1}^{9}a_{j}e^{0j}+\beta$, where $\beta$ is given by:
@@
Given a nilpotent matrix $M$, in the following theorem we give necessary and sufficient conditions on the tuple $\Lambda(M)$ to obtain pairwise non-isomorphic nilpotent almost abelian Lie algebras $\mathfrak{g}_{M}$ admitting symplectic structures. We exclude the zero matrix $M$, since when $M=0$, the Lie algebra $\mathfrak{g}_{M}$ is abelian and it admits a symplectic structure.
@@
Let $\mathfrak{g}_{M}={\mathbb{R}}e_{0}\ltimes_{M}{\mathbb{R}}^{2n-1}$ be the $2n$-dimensional nilpotent almost abelian Lie algebra with $\Lambda(M)=(n_{1},\ldots,n_{k};p_{1},\ldots,p_{k};t)$. Then $\mathfrak{g}_{M}$ admits a symplectic structure if and only if one of the following conditions holds:
@@ 2. ${\rm(ii)}$
We suppose that $\mathfrak{g}_{M}$ admits a symplectic structure. Hence, by Corollary 5.5, $M$ is conjugate to $C=\begin{pmatrix}0&0\\ w&D\end{pmatrix}$, where $w\in\mathbb{R}^{2(n-1)}$ and $D\in\mathfrak{sp}(2(n-1),{\mathbb{R}})$. Since $C$ is a non-zero nilpotent matrix, we have that, either $D=0$ and $w\neq 0$ or $D$ is a nonzero nilpotent matrix. If $D=0$, Corollary 3.7 implies that $C$, hence $M$, is conjugate to $\mathcal{J}_{2}\oplus 0_{2n-3}$. Therefore, $M$ satisfies condition (i). Now, we suppose that $D$ is a nonzero nilpotent matrix.
Then, Theorem 5.6 implies that $D$ is conjugate to
@@
Next we consider that $M$ is conjugate to the matrix given in the case (b). In this case $t$ is even, since $t=s$. We suppose that $m_{d}$ is even, so $m_{d}+1$ is odd and $m_{d-1}\geq m_{d}+1$. If $m_{d-1}>m_{d}+1$, then $\mathcal{J}_{m_{d}+1}$ appears only once. If $m_{d-1}=m_{d}+1$, then $m_{d-1}$ is odd. So Theorem 5.6 implies that $q_{d-1}$ is even. In this way, $\mathcal{J}_{m_{d}+1}$ appears an odd number of times. Now, we suppose that $m_{d}$ is odd. By Theorem 5.6, $q_{d}$ is even and Corollary 3.7 says that $\mathcal{J}_{m_{d}}$ appears $q_{d}-1$ times, that is, an odd number of times. Finally, the uniqueness in condition (ii) follows from Theorem 5.6.
@@
$t-1$ is even and $p_{i}$ is even for all $i$ such that $n_{i}$ is odd. Applying Theorem 5.6, $D$ is conjugate to a matrix in $\mathfrak{sp}(2(n-1),{\mathbb{R}})$. By Corollary 5.5, $\mathfrak{g}_{M}$ admits a symplectic structure.
@@
For each $n\in{\mathbb{N}}$, let
[…]
Consider the bijection $\tilde{\Lambda}$ defined in (8) for $b=0$ and $m=2n-1$. As a consequence of Lemma 3.3 and Theorem 5.10, we obtain the following classification result.
Corollary 5.11.
Let $n\in{\mathbb{N}}$. The isomorphism classes of $2n$-dimensional nilpotent almost abelian Lie algebras admitting a symplectic structure are parametrized by the set $\bigwedge^{s}_{2n-1}$.
The next corollary follows from the fact that $\bigwedge^{s}_{7}=\bigwedge_{7}$. This is in contrast with the complex case (see Example 4.12).
Corollary 5.12.
Any 8-dimensional nilpotent almost abelian Lie algebra admits a symplectic structure.
As a consequence of Theorem 4.10 and Theorem 5.10 we obtain the next result, whose proof follows from the straightforward fact that $\bigwedge^{c}_{2n-1}\subset\bigwedge^{s}_{2n-1}$.
@@
In Example 5.9 we showed by direct computation that the nilpotent almost abelian Lie algebra corresponding to the tuple $(5,3;1,1;1)$ admits no symplectic structure. It is easy to check that this is the only tuple in ${\bigwedge}_{9}-{\bigwedge}_{9}^{s}$. The isomorphism classes of 10-dimensional nilpotent almost abelian Lie algebras admitting neither complex nor symplectic structure are parametrized by the following intersection:
$\left({\bigwedge}_{9}-{\bigwedge}_{9}^{c}\right)\bigcap\left({\bigwedge}_{9}-{\bigwedge}_{9}^{s}\right)={\bigwedge}_{9}-{\bigwedge}_{9}^{s}=\{(5,3;1,1;1)\}.$
This implies the following corollary.
Corollary 5.14.
Up to isomorphism, the only $10$-dimensional nilpotent almost abelian Lie algebra admitting neither complex nor symplectic structure is $\mathfrak{g}_{M}$ with $\Lambda(M)=(5,3;1,1;1)$.
Remark 5.15.
We point out that if $\mathfrak{g}_{M}$ is a nilpotent almost abelian Lie algebra admitting no symplectic structure, then $\dim\mathfrak{g}_{M}\geq 10$. In fact, it follows from Theorem 5.8 that if $\mathfrak{g}_{M}$ is $k$-step nilpotent with $k\leq 4$, then $\mathfrak{g}_{M}$ admits a symplectic structure. Therefore, if $\mathfrak{g}_{M}$ admits no symplectic structure we must have $k\geq 5$. It was shown in [27] that the Lie algebra $\mathfrak{g}_{M}$ with $\Lambda(M)=(5;1;0)$ admits a symplectic structure. This forces the size of the matrix $M$ to be $m\times m$ with $m\geq 7$, that is, $\dim\mathfrak{g}_{M}\geq 8$. Since all $8$-dimensional nilpotent almost abelian Lie algebras admit a symplectic structure (see Corollary 5.12), it follows that $\dim\mathfrak{g}_{M}\geq 10$.
@@
Corollary 5.14 can be generalized by observing that, for $m_{1}$ and $m_{2}$ odd, with $m_{1}>m_{2}$, the tuple $(m_{1},m_{2};1,1;1)\in\bigwedge_{m\,}-\bigwedge_{m\,}^{s}$, for $m=m_{1}+m_{2}+1$. Therefore, the corresponding nilpotent almost abelian Lie algebra admits no symplectic structure. Moreover, by Theorem 5.13, this family of Lie algebras admits neither complex nor symplectic structure (see Corollary 5.21 for other examples).
Theorem 5.13 does not hold for arbitrary nilpotent Lie algebras, as the next example shows.
Example 5.17.